OpenStudy (anonymous):

What two numbers add up to -52 and are the product of -1260?

OpenStudy (anonymous):

let x be digit one and y be digit two

OpenStudy (anonymous):

x + y = -52 x * y = -1260

OpenStudy (anonymous):

y = -52 - x

OpenStudy (anonymous):

x * (-52-x) = -1260

OpenStudy (lovelycharm):

can u help me with mine

OpenStudy (anonymous):

Lovelycharm, sure.

OpenStudy (lovelycharm):

mine is 4x - 7y = 3 x - 7y = 27 @mjr632

OpenStudy (anonymous):

lovelycharm, isolate a single variable first.

OpenStudy (anonymous):

How could we isolate x in the second equation?

OpenStudy (lovelycharm):

we could isolate by cancel y and add 4 wish equal forsubtract the equal which equal 24 and divide 4 whish equal 6 righ @mjr632